OpenStudy (anonymous):

Fill in the missing term so that the quadratic equation has a graph that opens up, with a vertex of (– 2, – 16), and x intercepts at x = -6 and x = 2. (Do not include the negative sign in your answer.) y = x^2 + 4x − ___

OpenStudy (anonymous):

You would use the technique of completing the square for the vertex form of a parabola: y = a(x - h)^2 + k where (h, k) is the vertex Since the partial solution has an understood "1" as the coefficient on x^2, we have the simpler equation: y = (x - h)^2 + k -> y = [x - (-2)]^2 + (-16) y = (x + 2)^2 - 16 -> y = x^2 + 4x + 4 - 16 y = x^2 + 4x - 12

OpenStudy (anonymous):

All good now, @Melxo52 ?

OpenStudy (anonymous):

why thank you, you should be my math teacher! lol